OpenStudy (anonymous):

NEED HELP ASAP!!!!!!!!!!!!!!!!! Solve for n. ( 7 5/6 + 5 2/3 )+ 9 1/3= n + 15 A.7 1/3 B.7 5/6 C.6 2/3 D.6 5/6

OpenStudy (anonymous):

Plz help

OpenStudy (flower2468):

okay so find then common denominator or the bottom number of 3 and 6,

OpenStudy (anonymous):

You have to find the common denominator of the bottom numbers which is 3 and 6

OpenStudy (anonymous):

so the gcf of 3 and 6

OpenStudy (flower2468):

well, 6 is the common denominator so 6*1=6 and 3*2=6 and you have to do the same thing to the top numbers so 5*1=5 and 2*2=4 so the fractions with common denominators are 5/6 and 4/6, then you need to add them and that equals...?

OpenStudy (anonymous):

9/6

OpenStudy (anonymous):

yes

OpenStudy (flower2468):

correct and now add 7+5

OpenStudy (anonymous):

which equals 1 3/6 which is 1 1/2

OpenStudy (anonymous):

13 1/2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is that correct flower

OpenStudy (anonymous):

@sophissmart6330 which standard are you in?

OpenStudy (anonymous):

idk

OpenStudy (flower2468):

well the fraction is an improper fraction so if you 9-6=3 so there is another whle number so you add that to 13 and the fraction is 1/2 so ......14 1/2

OpenStudy (anonymous):

oh ok i was one off

OpenStudy (anonymous):

means of idk !

OpenStudy (anonymous):

sorry

OpenStudy (flower2468):

yes:) and now we need to add 14 1/2 and 9 1/3

OpenStudy (anonymous):

i am only in the sixth grade

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

23 5/6

OpenStudy (flower2468):

so we need to find the common denominator of 3 and 2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh, that's make. you need more practise on calculation of LCM

OpenStudy (anonymous):

1

OpenStudy (flower2468):

which is 6 so 3*2=6 and 2*3=6 and 1*2=2 and 1*3=3 so 2/6+3/6=5/6 and 9+14=23 so 23 5/6 is correct

OpenStudy (anonymous):

but it is not a answer choice

OpenStudy (flower2468):

so what +15=23 5/6

OpenStudy (anonymous):

oh ok

OpenStudy (flower2468):

23-15=....

OpenStudy (anonymous):

8

OpenStudy (flower2468):

so n=8 5/6

OpenStudy (anonymous):

look at the answer choices though

OpenStudy (anonymous):

there no 8

OpenStudy (anonymous):

plz help me

OpenStudy (flower2468):

oh okay then I guess I was wrong I will get someone I have to go @Lexi.babyy @-SpencerBraz-

OpenStudy (anonymous):

oh ok here is ur fan and medal

OpenStudy (flower2468):

thanks even though I don't deserve it

OpenStudy (anonymous):

u do

OpenStudy (anonymous):

ok plz help me anyone